Azobenzene (C6H5N)2 is produced from nitrobenzene (C6H5NO2) and triethyleneglycol (C6H14O4) in the presence of Zn and KOH according to the following balanced equation:



2 C6H5NO2 + 4 C6H14O4 → (C6H5N)2 + 4 C6H12O4 + 4 H2O


If 30.7 g of C6H5NO2 and 85.0 g of C6H14O4 are used and the reaction goes to completion, calculate the mass, in gram, of azobenzene produced by the reaction. Round your final answer to three significant figures.
1
Expert's answer
2019-09-30T05:21:58-0400

"n(C6H14O4)=85(g)\/150.17(g\/mol)=0.56 moles"

"n(C6H5NO2)=30.7(g)\/123(g\/mol)=0.25 moles"

Nitrobenzene is a limiting reagent.

"n((C6H5N)2)=n(C6H5NO2)\/2=0.25moles\/2=0.125moles"

"m(C6H5N)2)=182.2(g\/mol)*0.125moles=22.7g"

Answer: 22.7g